OpenStudy (anonymous):

Baby Anna weighed 96 ounces at birth and x ounces by the end of the first year. Baby Allison weighed 78 ounces at birth and gained twice as much weight as Baby Anna by the end of her first year. If at the end of the year both babies weighed the same amount, how much weight did Baby Allison gain?

OpenStudy (callisto):

Since ''Baby Allison weighed ... gained twice as much weight as Baby Anna by the end of her first year. '' Weight gained by Baby Alison = 2 times weight gained by baby Anna = 2x At the end of the year, weight of anna = 96 +x weight of alison = 78 +2x Since their weight are equal at the end of the year weight of anna =weight of alison 96 +x = 78 +2x Can you solve it from here? After you solve for x, Weight gained by alison = 2 times the value of x